OpenStudy (littleredriding):

Each small kite costs $4, while each large kite costs $7. If the total money collected selling kites is $150, and 30 kites are sold, how many of each size kite are sold? (Points : 3) 8 small and 22 large 10 small and 20 large 20 small and 10 large 27 small and 6 large

OpenStudy (littleredriding):

I say c. is that right?

OpenStudy (littleredriding):

@sammixboo @srgntb0$$ @Werewolfprincess

OpenStudy (littleredriding):

@srgntb0$$

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

ur right

OpenStudy (littleredriding):

Hi! Can you help me with this problem?

OpenStudy (littleredriding):

Oh okay thank you Logan!

OpenStudy (misty1212):

if you have choices, it is easy to check \[20+10=30\checkmark\] \[4\times 20+7\times 10=150\checkmark \]

OpenStudy (anonymous):

Solve:\[\{7 L+4 s=150,L+s=30\} \]

OpenStudy (littleredriding):

^ Thats what I did to get my answer

OpenStudy (anonymous):

Mathematica calculated {s = 20, L = 10}

OpenStudy (littleredriding):

Oh okay thank you!

OpenStudy (anonymous):

Your welcome.